OpenStudy (anonymous):

Find a polynomial P(x) with the conditions: Its degree is 4, its roots are 2, 3 and -1, and P(-2)=80.

OpenStudy (anonymous):

please answer, its an emergency

OpenStudy (anonymous):

okay since it's degree fora and we have three roots so we write it as follow: \[f(x)= a(x-2)(x-3))(x+1)^{2}\] of course I chose the term (x-1)^2 for example since the forth root is not stated,we could choose another root such as (x-1) for example ok now we find a:\[f(-2)=a(-4)*(-5)*(-1)^{2}\rightarrow 80=20a \rightarrow a=4\] so the f(x) that satisfies the conditions gives is : \[f(x)= 4(x-2)(x-3))(x+1)^{2}\] you can multiply now to expand it ..